alice, barbara and carol chip in for a $200 punch bowl for their mother. barbaras share is one and a half times alice's share and carol's share is $60. how much money did alice contribute toward the present?
Answer by TimothyLamb(4379)

a + b + c = 200
b = (3/2)a
c = 60
a + b = 200 - c
a + b = 200 - 60
a + b = 140
---
a + b = 140
a + (3/2)a = 140
(2/2)a + (3/2)a = 140
(5/2)a = 140
a = 140(2/5)
---
Answer:
a = $56
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php